# On Championing Foundation Models: From Explainability to Interpretability

### v1

@@ Abstract
Understanding the inner mechanisms of black-box foundation models (FMs) is essential yet challenging in artificial intelligence and its applications. Over the last decade, the long-running focus has been on their explainability, leading to the development of post-hoc explainable methods to rationalize the specific decisions already made by black-box FMs. However, these explainable methods have certain limitations in terms of faithfulness, detail capture and resource requirement. Consequently, in response to these issues, a new class of interpretable methods should be considered to unveil the underlying mechanisms in an accurate, comprehensive, heuristic and resource-light way. This survey aims to review interpretable methods that comply with the aforementioned principles and have been successfully applied to FMs. These methods are deeply rooted in machine learning theory, covering the analysis of generalization performance, expressive capability, and dynamic behavior. They provide a thorough interpretation of the entire workflow of FMs, ranging from the inference capability and training dynamics to their ethical implications. Ultimately, drawing upon these interpretations, this review identifies the next frontier research directions for FMs.
Keywords: Foundation models; explainable method; interpretable methods; machine learning theory
@@ 1 Introduction
Foundation Models (FMs), emerging as a new paradigm in deep learning, are fundamentally changing the landscape of artificial intelligence (AI). Unlike traditional models trained for specific tasks, FMs leverage massive and diverse datasets for training, which allows them to handle various downstream tasks through techniques like supervised fine-tuning (SFT) [1], Reinforcement Learning from Human Feedback (RLHF) [2], Retrieval Augmented Generation (RAG) [3], prompting engineering [4, 5] or continual learning [6]. In natural language processing, most FMs under autoregressive structure (e.g., GPT-3 [7], PaLM [8], or Chinchilla [9]) are established to output the next token given a sequence. Most text-to-image models, for example DALL·E [10], are trained to capture the distribution of images given a text input. Video FMs, trained on video data, potentially combined with text or audio, can be categorized based on their pretraining goals: generative models like VideoBERT [11] focus on creating new video content, while discriminative models like VideoCLIP [12] excel at recognizing and classifying video elements. Hybrid approaches like TVLT [13] combine these strengths for more versatile tasks. FMs have the potential to revolutionize a vast array of industries. Consider ChatGPT, a versatile chatbot developed by OpenAI, which has already made strides in enhancing customer support, virtual assistants, and voice-activated devices [14]. As FMs become more integrated into various products, their deployments will be scaled to accommodate a growing user base. This trend is evident in the growing list of companies, such as Microsoft with Bing Chat and Google with Bard, all planning to deploy similar FMs-powered products, further expanding the reach of this technology.
1.1 The Bottleneck of Explainability in Foundation Models
@@
Despite of rapid development of FMs, there are also ongoing concerns that continue to arise. The underlying causes behind certain emerging properties in FMs remain unclear. [15] firstly consider a focused definition of emergent abilities of FMs: “an ability is emergent if it is not present in smaller models but is present in larger models”. They conduct sufficient empirical studies to show emergent abilities can span a variety of large-scale language models, task types, and experimental scenarios. On the contrary, serveral works claim that emergent abilities only appear for specific metrics, not for model families on particular tasks, and that changing the metric causes the emergence phenomenon to disappear [16]. Moreover, prompting strategies can elicit chains of thought (CoT), suggesting a potential for reasoning abilities of FMs [5]. However, recent critiques by Arkoudas highlight limitations in current evaluation methods, arguing that GPT-4’s occasional displays of analytical prowess may not translate to true reasoning capabilities [17]. ‘All that glitters is not gold’, as potential risks gradually emerge behind the prosperous phenomenon. The phenomenon of “hallucination” has garnered widespread public attention, referring to the tendency of FMs to generate text that sounds convincing, even if it is fabricated or misleading, as discussed in recent research [18]. Additionally, studies have shown that adding a series of specific meaningless tokens can lead to security failures, resulting in the unlimited leakage of sensitive content [19]. These issues serve as stark reminders that we still lack a comprehensive understanding of the intricate workings of FMs, hindering our ability to fully unleash their potential while mitigating potential risks.
@@
However, explanations garnered from these explanable methods for complex FMs are often not entirely reliable and may even be misleading:
  • Inherent Unfaithfullness: [20] argue that if an explanation perfectly mirrored the model’s computations, the explanation itself would suffice, rendering the original model unnecessary. This highlights a key limitation of explainable methods for complex models: any explanation technique applied to a black-box model is likely to be an inaccurate representation. For instance, adversarial examples are carefully crafted inputs that are slightly perturbed from the original data, but often lead to misclassification or incorrect predictions by deep learning models. When using explainable methods to analyze and provide explanations for the model’s decisions on adversarial examples, the explanations may fail to capture the true underlying reasons for the misclassification.
  • Insufficient Detail: Even in a scenario where both the FMs and their explanation method are accurate – the models’ prediction is correct, and the explanation faithfully approximates it – the explanation might still be so devoid of detail as to be meaningless. For instance, consider an image of a dog with a frisbee in its mouth. An AI model can recognize it as a dog, but the explainable method might highlight only the dog’s overall shape or a few prominent features. The explanation may not reveal why the model made the correct prediction or the specific cues it used to recognize the dog.
  • Shallow Understanding: Current explainable methods primarily focus on addressing the crucial question of “how does the model achieve its output?", e.g., identifying which features play a significant role in decision-making. However, these methods often struggle to answer the fundamental questions that are particularly relevant to FMs, such as the underlying mechanisms that drive a model’s effectiveness, the power of scale, or the factors that trigger potential risks. These limitations stem from the inability of explanations to establish relationship between key factors such as model architecture, training data, and optimization algorithms.
  • Heavy Resource-dependency: As models become more complex, the difficulty of explaining them using traditional post-hoc methods like SHAP value [25], which often require significant computational resources, has increased significantly. This has limited the applicability of these techniques to large-scale models in real-world applications.
1.2 Interpretability tailored for Foundation Models
Numerous researchers have endeavored to distinguish the term “interpretability” from “explainability”. [40] define the interpretability of a machine learning system as “the ability to explain or present, in understandable terms to a human." [41] highlight, interpretability and explainability seek to answer different questions: ‘How does the model work?’ versus ‘What else can the model tell me?’. [20] states that interpretable machine learning emphasizes creating inherently interpretable models, whereas explainable methods aim to understand existing black-box models through post-hoc explanations. Following this definition, recent efforts have focused on developing inherently interpretable FMs based on symbolic or causal representations [42, 43, 44]. However, a prevalent reality in the field of FMs is that the most widely used and high-performing FMs remain black-box in nature.
@@ 1.2 Interpretability tailored for Foundation Models
From above all, there is no universally agreed-upon definition of interpretability within the machine learning community. The concept often varies based on the application domain and the intended audience [20, 45]. This motivates us to define appropriate interpretability and develop effective interpretable methods to reveal the internal workings of FMs, as shown in Figure 1. In our perspective, the interpretability tailored for FMs refers to the ability to reveal the underlying relationships between various related factors, such as model architecture, optimization process, data and model performance, in an accurate, legible, and heuristic way. This level of interpretability can significantly enhance public confidence in model components, mitigate potential risks, and drive continuous improvement of these black-box models. To achieve this, we will develop interpretable methods tailored for FMs, grounded in rigorous learning theory. These methods will encompass analyses of generalization performance, expressivity, and dynamic behavior. Generalization performance analysis, a form of uncertainty quantification, is the process of evaluating a foundation model’s ability to accurately predict or classify new, unseen data. Expressivity analysis is the process of evaluating a model’s capacity to represent complex patterns and functions within the data. Dynamic behavior analysis involves capturing the transformations and adaptations that occur within FMs as they learn from data, makes predictions, and undergoes training iterations. These interpretable methods align with the principles of interpretability as they:
  • Accurate. Compared to explainable methods based on post-hoc explanation, theoretical analysis, grounded in stringent mathematical or statistical principles, has the ability to predict the expected performance of algorithms under various conditions, ensuring accuracy and faithfulness in their assessments.
  • Comprehensive. The insights drawn from theoretical constructs are typically more comprehensive, delving into the multidimensional influences on model performance from a holistic viewpoint. For instance, within the realm of model generalization analysis, we can discern the intricate ways in which factors like optimization algorithms and model architecture shape and impact the model’s ability to generalize effectively.
@@
{forest}
[…]
Figure 2: The overall structure of the survey
@@
The structure of this paper is as follows (for a visual representation of the overall framework, please see Figure 2): Section 2 reviews key interpretable methods, including analyses of generalization performance, expressivity, and dynamic behavior. Section 3 delves into the inference capabilities of FMs, exploring in-context learning, CoT reasoning, and distribution adaptation. Section 4 provides detailed interpretations of FMs training dynamics. Section 5 focuses on the ethical implications of FMs, including privacy preservation, fairness, and hallucinations. Finally, Section 6 discusses future research directions based on the insights gained from the aforementioned interpretations.

### v2

# A Theoretical Survey on Foundation Models

@@ Abstract
Understanding the inner mechanisms of black-box foundation models (FMs) is essential yet challenging in artificial intelligence and its applications. Over the last decade, the long-running focus has been on their explainability, leading to the development of post-hoc explainable methods to rationalize the specific decisions already made by black-box FMs. However, these explainable methods have certain limitations in terms of faithfulness and resource requirement. Consequently, a new class of interpretable methods should be considered to unveil the underlying mechanisms of FMs in an accurate, comprehensive, heuristic, and resource-light way. This survey aims to review those interpretable methods that comply with the aforementioned principles and have been successfully applied to FMs. These methods are deeply rooted in machine learning theory, covering the analysis of generalization performance, expressive capability, and dynamic behavior. They provide a thorough interpretation of the entire workflow of FMs, ranging from the inference capability and training dynamics to their ethical implications. Ultimately, drawing upon these interpretations, this review identifies the next frontier research directions for FMs.
Keywords: Foundation models; explainable methods; interpretable methods; machine learning theory
@@ 1 Introduction
Foundation Models (FMs), emerging as a new paradigm in deep learning, are fundamentally changing the landscape of artificial intelligence (AI). Unlike traditional models trained for specific tasks, FMs leverage massive and diverse datasets for training, which allows them to handle various downstream tasks through approaches like supervised fine-tuning (SFT) [1], Reinforcement Learning from Human Feedback (RLHF) [2], Retrieval Augmented Generation (RAG) [3], prompting engineering [4, 5] or continual learning [6]. In natural language processing, most FMs under autoregressive structure (e.g., GPT-3 [7], PaLM [8], or Chinchilla [9]) are established to output the next token given a sequence. To learn the distribution of images conditioned on textual input, efforts have been made to propose text-to-image models, such as DALL·E [10]. Video FMs, trained on video data, potentially combined with text or audio, can be categorized based on their pretraining goals: generative models like VideoBERT [11] focus on creating new video content, while discriminative models like VideoCLIP [12] excel at recognizing and classifying video elements. Hybrid approaches like TVLT [13] combine these strengths for more versatile tasks. FMs have the potential to revolutionize a vast array of industries. Consider ChatGPT, a versatile chatbot developed by OpenAI, which has already made strides in enhancing customer support, virtual assistants, and voice-activated devices [14]. As FMs become more integrated into various products, their deployments will be scaled to accommodate a growing user base. This trend is evident in the growing list of companies, such as Microsoft with Bing Chat and Google with Bard, all planning to deploy similar FMs-powered products, further expanding the reach of this technology.
@@
Despite of rapid development of FMs, there are also ongoing concerns that continue to arise. The underlying causes behind certain emerging properties in FMs remain unclear. [15] firstly consider a focused definition of emergent abilities of FMs, “an ability is emergent if it is not present in smaller models but is present in larger models”. They conduct sufficient empirical studies to show emergent abilities can span a variety of large-scale language models, task types, and experimental scenarios. On the contrary, serveral works claim that emergent abilities only appear for specific metrics, not for model families on particular tasks, and that changing the metric causes the emergence phenomenon to disappear [16]. Moreover, prompting strategies can elicit chains of thought (CoT) for a reasoning task, suggesting a potential for reasoning abilities of FMs [5]. However, recent critiques by Arkoudas highlight limitations in current evaluation methods, arguing that GPT-4’s occasional displays of analytical prowess may not translate to true reasoning capabilities [17].
‘All that glitters is not gold’, as potential risks gradually emerge behind the prosperous phenomenon. The phenomenon of “hallucination” has garnered widespread public attention, referring to the tendency of FMs to generate text that sounds convincing, even if it is fabricated or misleading, as discussed in recent research [18]. Additionally, studies have shown that adding a series of specific meaningless tokens can lead to security failures, resulting in the unlimited leakage of sensitive content [19]. These issues serve as stark reminders that we still lack a comprehensive understanding of the intricate workings of FMs, hindering our ability to fully unleash their potential while mitigating potential risks.
1.1 Historical Notes on Explainability and Interpretability
@@
However, as discussed in Secton 2, these explainable methods have certain limitations in terms of faithfulness and resource requirement. Numerous researchers have endeavored to explore the term “interpretability”. [40] define the interpretability of a machine learning system as “the ability to explain or present, in understandable terms to a human.” [41] highlight, interpretability and explainability seek to answer different questions: ‘How does the model work?’ versus ‘What else can the model tell me?’. [20] states that interpretable machine learning emphasizes creating inherently interpretable models, whereas explainable methods aim to understand existing black-box models through post-hoc explanations. Following this definition, recent efforts have focused on developing inherently interpretable FMs based on symbolic or causal representations [42, 43, 44]. However, a prevalent reality in the field of FMs is that the most widely used and high-performing FMs remain black-box in nature.
@@
1.2 Objectives of This Survey
From above all, there is no universally agreed-upon definition of interpretability within the machine learning community. The concept often varies based on the application domain and the intended audience [20, 45]. This motivates us to define appropriate interpretability and develop effective interpretable methods to reveal the internal workings of FMs, as shown in Figure 1. In our perspective, the interpretability tailored for FMs refers to the ability to reveal the underlying relationships between various related factors, such as model architecture, optimization, data and model performance, in an accurate, comprehensive, heuristic and resource-light way. To achieve this, we will develop interpretable methods tailored for FMs, grounded in rigorous learning theory. These methods will encompass analyses of generalization performance, expressivity, and dynamic behavior. Generalization performance analysis, a form of uncertainty quantification, is an effective way to evaluate FMs’ ability to accurately predict or classify new, unseen data. Expressivity analysis is the process of evaluating a model’s capacity to represent complex patterns and functions within the data. Dynamic behavior analysis involves capturing the transformations and adaptations that occur within FMs as they learn from data, makes predictions, and undergoes training iterations. These interpretable methods align with the principles of interpretability as:
  • Accurate. Compared to explainable methods based on post-hoc explanation, theoretical analysis, grounded in solid mathematical or statistical principles, has the ability to predict the expected performance of algorithms under various conditions, ensuring accuracy in their assessments.
  • Comprehensive. The insights drawn from theoretical constructs are typically more comprehensive, delving into the multidimensional influences on model performance. For instance, within the realm of model generalization analysis, we can discern the intricate ways in which factors like optimization algorithms and model architecture shape and impact the model’s ability to generalize effectively.
@@ 1.2 Objectives of This Survey
Methods Generalization Expressive Power Dynamic Behavior
In-Context Learning -
Chain-of-Thought -
Adaption to Distribution Shifts - -
Emergent Capability - -
Parameter Evolution - -
RLHF - -
Privacy Leakage -
Unfair Prediction - -
Hallucination - -
Table 1: Various interpretable methods are employed to unveil distinct aspects of FMs. Generalization analysis is utilized to interpret In-Context Learning, Chain-of-Thought, Adaptation to Distribution Shift, Privacy Leakage, and Unfair Prediction. Expressive power analysis is applied to interpret In-Context Learning, Chain-of-Thought, and Hallucination. Dynamic behavior analysis is employed to interpret Emergent Capability, Parameter Evolution, RLHF, and Privacy Leakage.
@@
The structure of this survey is as follows: Section 2 shows the bottleneck of explainability in FMs. Section 3 reviews key interpretable methods, including analyses of generalization performance, expressivity, and dynamic behavior. The remaining sections aim to interpret the characteristics of FMs via three interpretable methods (as shown in Table 1). Section 4 delves into the inference capabilities of FMs, exploring in-context learning, CoT reasoning, and distribution adaptation. Section 5 provides detailed interpretations of FMs training dynamics. Section 6 focuses on the ethical implications of FMs, including privacy preservation, fairness, and hallucinations. Finally, Section 7 discusses future research directions based on the insights gained from the aforementioned interpretations.
2 The Bottleneck of Explainability in Foundation Models
Most explainable methods concentrates on providing users with post-hoc explanations on the specific decisions already made by the black-box models [20, 21, 22]. Despite their promising performance in explaining some black-box models, the explanations garnered from these methods for complex FMs are often not entirely reliable and may even be misleading.
2.1 Restricted Faithfulness
Faithfulness usually refers to how accurately the explainable methods reflect the true decision process of the black-box models [46]. Unfortunately, some studies indicate that these post-hoc explainable methods appear to be inherently unable to fully maintain faithfulness. [20] argue that if an explanation perfectly mirrored the model’s computations, the explanation itself would suffice, rendering the original model unnecessary. This highlights a key limitation of explainable methods for complex models: any explanation technique applied to a black-box model is likely to be an inaccurate representation. Based on that fact, the mainstream work focuses on evaluating the faithfulness of post-hoc explainable methods from different perspectives, such as consistency and sufficiency [47], usefulness to humans [48, 49] and robustness [50].
Figure 2: Heatmap displaying attention weights generated from a negative movie review. On the left, we present the model’s actual attention pattern on, while on the right, we showcase a set of attention weights created adversarially. Despite their significant differences, both patterns result in the same prediction (0.01) effectively.
From our perspective, comprehensiveness appears as an important metric for evaluating the faithfulness of FMs. Comprehensiveness specifically denotes that the explanations provided should be closely linked to the FMs’ performance. Taking an example to illustrate the limitation of explainable methods in terms of comprehensiveness. Due to the capability to capture the long-distance dependency relationship between features, self-attention mechanisms have been widely-used in various FMs. The self-attention visualized by some explainable methods such as BertViz [51], is commonly viewed as a explanation for the working mechanism of FMs by presenting the relative importance of units. However, we may drawn misleading relationship between theses attention weights and model outputs (as shown in Figure 2 shown in [52]). The left panel displays the initial attention distribution $\alpha$ across the words in a specific movie review using a conventional attentive BiLSTM setup for sentiment analysis. It might be tempting to infer from this that the token “waste” predominantly influences the model’s ‘negative’ classification ($\hat{y}=0.01$). However, an alternative attention distribution $\tilde{\alpha}$ (right panel) could be crafted to focus on entirely different tokens while still producing a nearly identical prediction (assuming all other parameters of $f$, $\theta$, remain constant). This means that the discovered attention is linked to the model’s performance at a lower level.
2.2 Heavy Resource-dependency
As models become more complex, the difficulty of explaining them using traditional post-hoc methods like SHAP [25] and Local Interpretable Model-agnostic Explanations (LIME) [24]. Despite the valuable explanations provided by these methods, they have limitations in terms of computational complexity. For example, for each testing sample, the weight parameters derived from local linear models are integrated into SHAP. Nevertheless, SHAP necessitates models to be constructed for every testing or validation sample, leading to a computationally expensive process when globally calculating SHAP values.
TreeSHAP Version Time Complexity Space Complexity
Original TreeSHAP $O(MTLD^{2})$ $O(D^{2}+|N|)$
Fast TreeSHAP v1 $O(MTLD^{2})$ $O(D^{2}+|N|)$
Fast TreeSHAP v2 (General case) $O(TL2^{D}D+MTLD)$ $O(L2^{D})$
Fast TreeSHAP v2 (Balanced case) $O(TL^{2}+MTLD)$ $O(L^{2})$
Table 2: Summary of space complexity and computational complexities of TreeSHAP algorithms [53]
Recently, there are two ways to mitigate the highly computational complexity with SHAP. Firstly, to address the computational challenges, a novel class of stochastic estimators is proposed by leveraging feature subset or permutation sampling [54, 55]. While these estimators provide consistent results, their high computational complexity, often requiring many model evaluations, hinders their scalability and real-world deployment. Secondly, previous work has explored model-specific Shapley value approximations for trees [56] and neural networks [57, 58, 59], often leveraging their unique structures. While these methods can reduce computational cost, they may introduce bias in feature attribution and struggle to accurately handle features that were not seen during training [60, 61, 62]. Additionally, these methods result in a limited performance in computational cost when explaining large models with huge amount data. Table 2, borrowed from [53], presents the summary of computational complexities of TreeSHAP algorithms, where $D$ is the maximum depth of a tree, $|N|$ stands for the dimension of features, $T$ denotes the number of trees, $M$ represents the total samples size, and $L$ indicates the maximum number of leaves in a tree. From this table, even the accelerated variant of TreeSHAP reveals a considerable computational overhead when elucidating complex models with substantial datasets.
@@ 3 Interpretable Methods
3.1 Preliminary
Before the analysis of generalization, expressive power, and dynamic behavior, we first introduce the preliminary of the target of FMs and notions for the following analysis.
FMs represent a class of machine learning techniques that utilize large-scale neural networks, characterized by multiple layers of neurons and nonlinear activation functions. These models are engineered to process and learn from diverse training datasets, extracting intricate patterns and features through a series of computational layers. The data they utilize can come from real-world records, human interactions, or even synthetic data generated by other models. By leveraging proper architectures and training algorithms, empirical results show that FMs can autonomously uncover and represent intricate relationships within most types of data. This enables them not only to make sophisticated predictions and informed decisions but also to generate high-quality synthetic data based on given prompts and the extensive knowledge they have accumulated.
In real-world application, FMs serve the purpose of providing reasonable, meaningful, and helpful responses to user inputs. Mathematically, the goal of FMs is to approximate a mapping that aligns with what humans would consider an ideal output. This mapping treats $\mathcal{X}$ and $\mathcal{Y}$ as the input and output spaces, respectively. To achieve this, each FM is accessible to a set of input-output pairs
$\mathcal{S}=\{(x,y):x\in\mathcal{X},y\in\mathcal{Y}\}.$
where $x$ is the input and $y$ is one of the reasonable output. For example, consider a language translation FM designed to translate English sentences into French. Here, $\mathcal{X}$ would be the space of all possible English sentences, and $\mathcal{Y}$ would be the space of all possible French sentences. The set $\mathcal{S}$ would consist of pairs of English sentences and their corresponding French translations, such as (“Hello, how are you?", “Bonjour, comment ça va?").
To achieve better performance, the accessible dataset $\mathcal{S}$ is always divided into three subsets: training, validation, and testing sets. The training set is used to teach the FM by allowing it to learn the intricate patterns and relationships within the data. This subset forms the core of the learning process, enabling the model to adjust its parameters and improve its performance iteratively. The validation set, on the other hand, is used to evaluate the model’s performance during the training process. It helps in selecting the best model architecture and hyperparameters by providing a separate dataset that the model has not seen during training. This ensures that the model generalizes well to new, unseen data and prevents overfitting to the training set. Finally, the testing set is used to evaluate the model’s performance objectively. This subset is completely isolated from the training and validation processes, providing an unbiased assessment of how well the FM can handle real-world data.
It is important to note that not all $(x,y)$ pairs are ideal; some pairs may be noisy or contain errors, reflecting real-world imperfections. Additionally, in some cases, the output $y$ may not be provided, corresponding to unsupervised learning scenarios.
Remark 3.1.
It’s crucial to acknowledge that we always assume samples in $\mathcal{S}$ is i.i.d, where each data point is assumed to be drawn from the same distribution $\mathcal{D}$ and is unrelated to other data points, may not hold in many real-world scenarios. For instance, in recommendation systems, user interactions and preferences are often influenced by various factors such as temporal dynamics, social influences, and contextual information, leading to data that is inherently non-i.i.d. Despite these complexities and the non-i.i.d nature of real-world data, FMs have shown remarkable robustness and effectiveness. They are capable of capturing intricate patterns and interactions within the data, thereby maintaining strong performance even when the i.i.d assumption is violated. However, for the purpose of this survey, we will primarily focus on the scenarios where the i.i.d assumption is maintained.
With the accessible dataset, FMs first map all $x\in\mathcal{X}$ to $R^{d_{X}}$ and $y\in\mathcal{Y}$ to $R^{d_{Y}}$. This transformation is crucial in various applications as it allows models to work with structured data representations. For example, in image classification tasks, $(x,y)$ represents an image-label pair, where the image is mapped to a vector of pixel values and the label is transformed into a one-hot encoded vector. In this survey, this process is not important thus we just let $\mathcal{X}=R^{d_{X}}$ and $\mathcal{Y}=R^{d_{Y}}$.
Then, FMs are trained to learn a hypothesis, denoted by $h$, from the training data. Due to the specified architecture and numerical precision established by FMs, there are constraints on $h$. Mathematically, all such possible hypotheses constitute the hypothesis space, denoted by $\mathcal{H}$. For instance, when an FM is a neural network, $\mathcal{H}$ is denoted as
$\{h:h(x)=W_{D}\sigma_{D-1}(W_{D-1}\sigma_{D-2}(\ldots\sigma_{1}(W_{1}x)))\}$
where the depth of FMs, the $j$-th weight matrix and the $j$-th activation function are denoted by $D$, $W_{j}$ and $\sigma_{j}$. Here, the activation function $\sigma_{j}$ is assumed to be continuous such that $\sigma_{j}(0)=0$. Popular activation functions include the softmax, sigmoid, and tanh functions.
To interpret FMs, theoretical research can be categorized into 3 parts: generalization analysis, expressive power analysis, and dynamic behavior analysis. Their purpose and motivation are as follows:
  1. 1.
    Generalization analysis: This analysis focuses on understanding how well foundation models generalize from the training data to unseen data. The goal is to theoretically explain the factors influencing the generalization performance, such as model complexity, data distribution, and the training dynamics. It provides insights into why FMs can perform well in diverse applications even with limited training examples in some domains.
  2. 2.
    Expressive power analysis: This part examines the capacity of foundation models to represent complex functions or tasks. The motivation is to quantify the types of functions that the model can approximate, given its architecture. Understanding the expressive power helps in determining the potential and limitations of FMs in capturing the intrinsic structure of different data types, thus guiding architectural improvements.
  3. 3.
    Dynamic behavior analysis: This analysis investigates the training dynamics of FMs, including optimization trajectories and the evolution of model parameters during training. The purpose is to understand how the model’s behavior changes during training and how different training techniques, such as learning rate schedules or regularization, impact the final model. This perspective is crucial for improving training stability and efficiency, as well as for mitigating issues like catastrophic forgetting or mode collapse.